OpenStudy (anonymous):

Help with calculus. Find the dimensions of the right circular cone with maximun volume that is possible to inscribe on a sphere of radius r. I know that i have to maximize ( using derivatives ) the volume, but im still figuring out the relationship between the volume of the cone and the radius of the sphere. Thanks

ganeshie8 (ganeshie8):

sphere : x^2 + y^2 + z^2 = r^2

OpenStudy (loser66):



OpenStudy (loser66):

the volume of the cone w.r.t the height and the radius of the base. compare the 2, which one is max value of the cone?

OpenStudy (loser66):

don't forget \(V_{cone}=\pi r^2 *h\) ( I mean r is square)

OpenStudy (loser66):

and r is maximum when the base of the cone at the middle of the sphere. ha!! I don't know how to say.

OpenStudy (loser66):

@ganeshie8 help me, please. I don't know how to explain